OpenStudy (anonymous):

I am at my wits end with this problem. Solve the system of equations. 2x + 4y + 3z = 6 3x + 5y + 6z = 3 2x + 3y + 4z = 8 The answer choice are a. (x= 63, y= -18, z= -16) b. (x= 65, y= -20, z= -14) c. (x= 62, y= -17, z= -15) d. (x= 64, y= -19, z= -17)

OpenStudy (anonymous):

i think it's A.

OpenStudy (anonymous):

hold on don't do it yet okay

OpenStudy (anonymous):

Okay I'll wait

OpenStudy (jdoe0001):

you take 2 equations at a time, then cancel out 1 of the variables, in the end, you'll end up with 2 variables system of equations 2x + 4y + 3z = 6 <---- let's use this one 3x + 5y + 6z = 3 <----- and this one 2x + 3y + 4z = 8 let's cancel out by elimination 2x + 4y + 3z = 6 let's multiply this one by - 2 3x + 5y + 6z = 3 -4x - 8y - 6z = -12 3x + 5y + 6z = 3 -------------------- -x -3y + 0 = -9 <--- 1st 2 variable one let's do the another 2 2x + 4y + 3z = 6 <--- let's use this one again 3x + 5y + 6z = 3 2x + 3y + 4z = 8 <--- and this one 2x + 4y + 3z = 6 <----- let's multiply this one by 4 2x + 3y + 4z = 8 <---- let's multiply this one by - 3 8x + 16y + 12z = 24 -6x - 9y - 12z = -24 ------------------------- 2x + 7y + 0 = 0 <---- 2nd 2 variable one

OpenStudy (anonymous):

Yes but I need the solutions. I can get it down to where I eliminate variables but none of my answers make sense.

OpenStudy (jdoe0001):

\(\bf -x -3y + 0 = -9 \\ 2x + 7y + 0 = 0\\\quad \\\quad \\ -x-3y=-9\\ 2x+7y = 0 \) then solve it like you'd any 2 variable system of equations

OpenStudy (jdoe0001):

woops, need to dash

OpenStudy (anonymous):

Thank you sooo much!

OpenStudy (anonymous):

i am still here i will try

OpenStudy (anonymous):

dam it this is hard i will still try it's not A

OpenStudy (anonymous):

Idk it does not work i tryed i did my best i could not find it

OpenStudy (anonymous):

i am sorry about this when S17 gets on he would be able to help @S17

OpenStudy (anonymous):

sorry clan member i could not find it ^---^

OpenStudy (anonymous):

@campbell_st help would you please we can not find it out

OpenStudy (campbell_st):

well I wouldn't solve anything, I'd just substitute solution 1 x = 63, y = -18, z = -16 substitute into the 1st equation 2(63) + 4(-18) + 3(-16) = 6 true so test this on the 2nd equation 3x + 5y+ 6z = 3 3(63) + 5(-18) + 6(-16) = 3 true so try it on the 3rd equation 2x + 3y + 4z = 8 2(63) + 3(-18) + 4(-16) = 8 so the values x = 63, y = -18 and z = -16 is a solution to the system of equations.

OpenStudy (anonymous):

did he help at all

OpenStudy (anonymous):

so it would be A.